OpenStudy (anonymous):

Prove that if m and n are integers and mn is even, then m is even or n is even. Use the contrapositive (indirect proof).

OpenStudy (freckles):

Have you found the contrapositive?

OpenStudy (anonymous):

No

OpenStudy (freckles):

We are trying to write the contrapositive of: For all integers m and n, if mn is even, then m is even or n is even. Instead of writing if p then q we need to write if not q then not p. So this means we to negate m is even or n is even We also need to negate mn is even --- hint for negating that or statement: not(a or b)=not a and not b

OpenStudy (anonymous):

n is odd or m is odd mn is odd?

OpenStudy (freckles):

look at the hint for negating an or statement you need to change or to and

OpenStudy (anonymous):

yeah I just saw that so n is odd and m is odd?

OpenStudy (freckles):

So you need to prove: If n and m are odd integers, then mn is odd.

OpenStudy (anonymous):

n = 2x and m = (2y +1) n + m = (2x + (2y +1)) = 2(x + (y + 1))

OpenStudy (freckles):

since n is odd then you need to represent n as a odd integer not an even integer so n=2x+1 and m=2y+1

OpenStudy (freckles):

and also you are asked to find nm not n+m

OpenStudy (anonymous):

umm im thinking

OpenStudy (freckles):

nm means n times m

OpenStudy (anonymous):

n=2x+1 and m=2y+1 nm = (2x +1)(2x +1)

OpenStudy (freckles):

ok then multiply the (2x+1) and (2y+1) I assume you meant nm=(2x+1)(2y+1)

OpenStudy (anonymous):

I get 4xy +2x +2y + 1. Not sure if that is odd

OpenStudy (freckles):

see if you can put an the odd looking form which is 2*some integer+1 or 2k+1 where k is some integer

OpenStudy (freckles):

look at those first three terms you wrote

OpenStudy (freckles):

all of them have a factor 2 in common

OpenStudy (anonymous):

which terms? I am sorry I am lost

OpenStudy (freckles):

4xy,2x,2y

OpenStudy (freckles):

those first three terms of 4xy+2x+2y+1

OpenStudy (freckles):

they all have a 2 in common (the first terms do)

OpenStudy (freckles):

three terms*

OpenStudy (anonymous):

so how do you use 2k + 1?

OpenStudy (freckles):

you want to write 4xy+2x+2y+1 in the form that is 2k+1

OpenStudy (anonymous):

I am not sure how to do that?

OpenStudy (freckles):

Do you know how to factor?

OpenStudy (freckles):

There is a factor 2 in the first three terms

OpenStudy (anonymous):

2(2x^2 + y) + 1?

OpenStudy (freckles):

4xy+2x+2y+1 2*2xy+2*x+2*y+1 2(2xy+x+y)+1 Isn't this in the form 2k+1 2xy+x+y is an integer?

OpenStudy (freckles):

Not sure how you got the square

OpenStudy (anonymous):

I am sorry Im a little lost

OpenStudy (freckles):

4xy,2x,and 2y all have a common factor 2 so I factored a 2 from all three them in the expression 4xy+2x+2y+1 2(2xy+x+y)+1

OpenStudy (anonymous):

ok so that is in form of 2k +1?

OpenStudy (anonymous):

k = 2xy + x + y?

OpenStudy (freckles):

\[4xy+2x+2y+1 \\ 2 \cdot 2xy+ 2 \cdot x +2 \cdot y +1 \\ 2(2xy+x+y)+1\] Do you think it is in the form 2k+1?

OpenStudy (freckles):

Is 2xy+x+y an integer?

OpenStudy (freckles):

Is the set of integers closed under addition and multiplication?

OpenStudy (freckles):

x,y are integers

OpenStudy (freckles):

When you mutliply x and y are multiply them with any other integer you will get an integer and when you add them with any other integer you will get an integer 2xy+x+y is an integer 2(2xy+x+y)+1 is in the form 2k+1

OpenStudy (freckles):

Therefore mn is odd

OpenStudy (anonymous):

Oh ok so you plug (2xy + x + y) into k?

OpenStudy (freckles):

You don't have to plug into k You are just trying to show you can write mn in the form that is 2k+1 We did that by writing mn as 2(2xy+x+y)+1

OpenStudy (anonymous):

oh I see it. Thank you very much!

OpenStudy (freckles):

np

OpenStudy (loser66):

Suppose neither of them are even, that means both m, n are odd, and mn is even so that m, n can be written as m = 2k +1 n = 2s +1 for some k, s in Z mn = (2k+1)(2s +1) = 4ks +2k+2s +1= 2(2ks +k +s) +1, let 2ks +k +s = l, then mn = 2l +1 form the form of an odd number (contradict) --> It is not the case that both them are odd or one of them must be even